OpenStudy (anonymous):

a rectangle park is 12 km long and 8 km wide. how long is a pedestrain route that runs diagoinally across the park?

jimthompson5910 (jim_thompson5910):

solve for c to get your answer a^2 + b^2 = c^2 12^2 + 8^2 = c^2 144 + 64 = c^2 ... ... ... c = ???

OpenStudy (anonymous):

than the answer would be 50 right?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (whpalmer4):

144 + 64 = ?

OpenStudy (anonymous):

I meat 4 sqrt 13km

jimthompson5910 (jim_thompson5910):

\[\large 4\sqrt{13}\] is correct

OpenStudy (whpalmer4):

Numerically, \(4\sqrt{13}\approx 14.42\)